OpenStudy (anonymous):

if a-b=10 and a^2-b^2=20, what is the value of b?

OpenStudy (anonymous):

divide first equation into second gives a + b = 2 solving sumultaneously a-b=10 a+b = 2 add 2a=12 a=6 so b =2-6 = -4

OpenStudy (anonymous):

sorry i should of explained the division : a^2-b^2 = (a-b)(a+b) (difference of 2 squares) so a^2-b^2/(a-b) = (a-b)(a+b)/(a-b) = a+b and of course 20/10 = 2